Suppose Blue Thumb Tools is considering the introduction of a new, heavier hammer to be used...

  1. Suppose Blue Thumb Tools is considering the introduction of a new, heavier hammer to be used for driving spikes. The new hammer will cost $490,000. The cost will be depreciated straight-line to zero over the project’s five-year life, at the end of which the new hammer can be scrapped for $40,000. The new hammer will save the firm $146,000 per year in pretax operating costs, and it required an initial investment in net working capital of $35,000. The tax rate of the firm is 30%. What are the cash flows of firm’s new project (using a time line)?
Answer #1

1)

Cash flow at year 0 = Initial investment + NWC

Cash flow at year 0 = -490,000 - 35,000

Cash flow at year 0 = -525,000

2)

Deprecation = 490,000 / 5 = 98,000

Cash flow at year 1 = (Pre tax savings - depreciation)(1 - tax) + depreciation

Cash flow at year 1 = (146,000 - 98,000)(1 - 0.3) + 98,000

Cash flow at year 1 = 33,600 + 98,000

Cash flow at year 1 = $131,600

3)

Cash flow at year 2 = $131,600

4)

Cash flow at year 3 = $131,600

5)

Cash flow at year 4 = $131,600

6)

No operating cash flow at year 5 = Market value + NWC - tax(market value - book value)

No operating cash flow at year 5 = 40,000 + 35,000 - 0.3(40,000 - 0)

No operating cash flow at year 5 = 40,000 + 35,000 - 12,000

No operating cash flow at year 5 = 63,000

Cash flow at year 5 = $131,600 + $63,000

Cash flow at year 5 = $194,600